Solid metal cube of side 10cm is heated from 10°c. if the linear expansivity of the metal is 1.2×10-5k-1, calculate the increase in it volume

1
Expert's answer
2021-12-09T09:22:40-0500

We can find the increase in the volume of the metal cube as follows:


"\Delta V=3\alpha V\Delta T,""\Delta V=3\times1.2\times10^{-5}\ (^{\circ}C)^{-1}\times(10\ cm)^3\times(60^{\circ}-10^{\circ}),""\Delta V=1.8\ cm^3."